OpenStudy (kendricklamar2014):

Simplify 3! A. 2 B. 5 C. 3 D. 6

OpenStudy (a1234):

3*2*1 = 6

OpenStudy (anonymous):

^^^^^

OpenStudy (anonymous):

yeah! 3! = 3*2*1 = 6 :D

OpenStudy (anonymous):

Fun Factorials XD